# Efficient and high-performance routing of lattice-surgery paths on three-dimensional lattice

### v3

@@ Abstract
Encoding logical qubits with surface codes and performing multi-qubit logical operations with lattice surgery is one of the most promising approaches to demonstrate fault-tolerant quantum computing. Thus, a method to efficiently schedule a sequence of lattice-surgery operations is vital for high-performance fault-tolerant quantum computing. A possible strategy to improve the throughput of lattice-surgery operations is splitting a large instruction into several small instructions such as Bell state preparation and measurements and executing a part of them in advance. However, scheduling methods to fully utilize this idea have yet to be explored. In this paper, we propose a fast and high-performance scheduling algorithm for lattice-surgery instructions leveraging this strategy. We achieved this by converting the scheduling problem of lattice-surgery instructions to a graph problem of embedding 3D paths into a 3D lattice, which enables us to explore efficient scheduling by solving path search problems in the 3D lattice. Based on this reduction, we propose a method to solve the path-finding problems, look-ahead Dijkstra projection. We numerically show that this method reduced the execution time of benchmark programs generated from quantum phase estimation algorithms by 3.8 times compared with a naive method based on greedy algorithms. Our study establishes the relation between the lattice-surgery scheduling and graph search problems, which leads to further theoretical analysis on compiler optimization of fault-tolerant quantum computing.
@@ 1 Introduction
Quantum computing can efficiently solve several vital scientific problems [shor1999polynomial, kitaev1995quantum, lloyd1996universal], but its high error rates prevent us from a practical demonstration of quantum advantage. We can overcome this problem with a quantum error-correction technology, i.e., by creating a logical qubit with a negligible error rate by using several noisy qubits. Surface codes [kitaev1997quantum, bravyi1998quantum, fowler2012surface] are known as promising candidates for quantum error-correcting codes since they show high error-correction performance and can be implemented with qubits having nearest-neighbor interactions on a two-dimensional (2D) grid lattice. Lattice surgery [horsman2012surface, fowler2018low] plays a key role in the fault-tolerant quantum computation with surface-code logical qubits: they allow multi-qubit Pauli measurements in a fault-tolerant manner by connecting target logical qubits through ancillary logical-qubit space, and they can be performed in parallel as long as their paths do not overlap. The current state-of-the-art resource estimation stands on architectures with surface codes and lattice surgery [beverland2022assessing, babbush2018encoding, yoshioka2022hunting].
@@
To run quantum algorithms with a short execution time, we should maximize the throughput of quantum instructions [javadiabhari2014scaffcc, Qiskit, sivarajah2020t, watkins2023high] by optimizing their scheduling. Unfortunately, finding the optimal scheduling of lattice-surgery instructions is known to be NP-hard [herr2017optimization, molavi2023compilation]. Thus, fast and near-optimal approximation strategies for scheduling lattice-surgery instructions are demanded. For frequently used modules such as magic-state distillation circuits, there are dedicated designs and logical-qubit mappings [fowler2018low, litinski2019game, litinski2019magic]. On the other hand, for general purposes, we need an algorithm that can accept an arbitrary lattice-surgery sequence. One possible strategy is to map problems into well-known NP-hard instances and find high-quality solutions with fast solvers. Lao et al. [lao2018mapping] map problems into the quadratic assignment problem, and Molavi et al. [molavi2023compilation] into SAT problems. While this method would show near-optimal solutions, the target size of the compilation is limited by the capability of solvers. Thus, efficient and high-performance scheduling methods based on other concepts are still in demand.
@@
One of the promising approaches for general and high-performance optimization is to split instructions that are difficult to parallelize into several small fragments and execute parts of them in advance. Beverland et al. [beverland2022surface] proposed a method based on this idea. They propose methods to find a way to split and schedule instructions by finding the edge-disjoint path on a 2D lattice and converting the solution to a lattice-surgery sequence for multiple execution cycles. While this strategy is shown to be effective, the proposed method sticks to specific logical-qubit allocation patterns while possible logical-qubit allocations can be arbitrary [babbush2018encoding, beverland2022assessing, yoshioka2022hunting], which loses the versatility in compilation targets. Also, the range of optimization in this algorithm is limited to instructions within a narrow time window. Therefore, we need a strategy that can fully leverage the idea of instruction split and preceding execution of them.
@@
Here, we propose an efficient and high-performance lattice-surgery scheduling algorithm that utilizes the strategy. The key step of our proposal is establishing the correspondence between a path in a 3D lattice, an executable sequence of lattice-surgery instructions, and achievable actions on target logical qubits. Thanks to this correspondence, we can rephrase the optimization of a lattice-surgery sequence to a problem to embed several 3D paths in a 3D grid lattice, in which we can utilize graph optimization methods. On the basis of this result, we show several polynomial-time approximation algorithms to solve this problem. We evaluated these algorithms with numerical simulations and found that one among them, named look-ahead Dijkstra projection, significantly improves throughput with a reasonable compilation time. We numerically showed that the look-ahead Dijkstra projection improves throughput by 3.8 times compared with a naive solution in a benchmark with realistic applications, SELECT modules in qubitization-based quantum phase estimation [low2019hamiltonian, babbush2018encoding].
@@
Our results not only optimize the runtime of quantum algorithms but also show the complicated lattice-surgery compilation can be reduced to simple graph problems, where we can naturally utilize the knowledge of path-search algorithms. Since our algorithm works for general logical-qubit mapping and connectivity, our methods can be extended to other fault-tolerant quantum computing (FTQC) designs based on lattice surgery or distributed quantum computing. While our benchmark focuses on the case of two-body lattice-surgery instructions, we also showed that our method can be straightforwardly extended to multi-body lattice surgery in Appendix. A. Thus, our results offer a base for versatile and high-performance compilation for fault-tolerant quantum computing.
@@
  • We propose an efficient and high-performance scheduling method for our 3D formulation, called the look-ahead Dijkstra projection. This algorithm leverages the properties of the extended formalism with only a modest increase in time complexity compared to simple baseline algorithms.
@@ 2.1 Surface codes and logical operations
Since qubits suffer from large error rates, we need quantum error correction to suppress its error rate to an arbitrarily small value. Surface codes are known as the most promising quantum error-correcting codes, which encode the information of a qubit into a square cell of qubits. The width of the cell corresponds to a code distance, and by increasing the code distance, we can exponentially suppress the error rates if physical error rates are smaller than the value known as thresholds [kitaev1997quantum, fowler2012towards, acharya2022suppressing]. A cell of surface codes has two types of boundaries: $X$-type and $Z$-type, where Pauli-$X$ operations along with $X$-type boundary constitute logical $X$ operations, and the same for Pauli-$Z$.
We suppose that qubits are integrated on a large two-dimensional plane, which we call a qubit plane. We assume that the qubit plane is divided into several surface-code cells. During the computation, several cells are used for keeping single-qubit information, which we call data cells. The other cells are available as working space for logical operations on the data cells, which we call ancillary cells. Fig. 1 shows an example of a qubit plane, where dark cells are data cells and light cells are ancillary cells. As surface codes have the two types of boundaries, each cell also has two types of directions: top and bottom are $X$-boundaries and the others are $Z$-boundaries, or top and bottom are $Z$-boundaries and the others are $X$-boundaries.
@@ 2.2 Scheduling of lattice surgery
Among the universal operations, the scheduling of lattice-surgery instructions tends to be a complicated task since multiple cells are relevant to them. Thus, in this paper, we focus on the scheduling of lattice-surgery instructions. A lattice-surgery instruction is achieved by connecting target logical-qubit cells. If the target cells are not neighboring, we can use ancillary cells to connect them. Once connected, we need to repeat syndrome measurements for $d$ times, where $d$ is a code distance, to guarantee reliable parity-check measurements. Thus, it is convenient to define the duration for lattice surgery as a time unit, which is called code beat. The types of connected boundaries determine the basis of Pauli measurements. We can perform Pauli-$Z$ measurements if a $Z$-boundary of a cell is connected and Pauli-$X$ for $X$-boundary.
@@
This example illustrates that optimizing the ordering and routing of lattice surgery is an important factor in maximizing the performance of quantum computing. Unfortunately, the optimal scheduling of the sequence of lattice surgery is known to be NP-hard, i.e., the optimal scheduling is computationally difficult to find within reasonable time [herr2017optimization]. Thus, an efficient algorithm that can find the near-optimal paths of lattice-surgery instructions is strongly demanded.
@@ 2.3 Baseline scheduling method: BFS
As a baseline method, we introduce a simple procedure to schedule the sequence of lattice-surgery instructions. The most naive strategy for scheduling instructions is to use breadth-first search (BFS) as a subroutine to find one of the shortest paths. We call the procedure as BFS, which is defined as follows.
@@
The pseudo-code of this procedure is shown in Alg. 1. Here, we keep whether the ancillary cells are available or not in a map object $A$. FindShortestPath will return one of the shortest paths from $v_{1}$ to $v_{2}$ on graph $G$ for a lattice surgery instruction $o=(v_{1},v_{2},d)$ so that the path does not use surface-code cells marked $1$ in $A$ and $v_{1},v_{2}$ are connected to the path from $d\in\{X,Z\}$ boundaries. If there is no path between them, the function returns an empty list. We can find the shortest paths in a graph $G$ with $V$ vertices and $E$ edges in $O(V+E)$ time using BFS. In this procedure, while choosing the shortest path will not necessarily lead to optimal results, we expect this choice to be reasonable since it will consume the fewest cells.
@@ 3.2 Motivating examples
Here, with a simple example, we show that splitting of lattice-surgery instructions enables efficient use of resources, i.e., voxels in the 3D lattice. An overview of our example is shown in Fig. 4. We consider the same situation as Fig. 2. In this situation, BFS or look-ahead BFS algorithms cannot schedule all the routing at the same code beat, and a hatched path between green cells must wait for the next code beat. The 3D-lattice representation of this scheduling is shown in the left half of the figure. In this case, voxels below green paths are left unused and lose the parallelism of scheduling.
@@
If there is a 3D path that connects XX-($Z$-)boundaries of cells and has an even number of kinks, then there is a sequence of lattice-surgery instructions that consumes the resource corresponding to voxels in the path and results in logical Pauli-XXXX(-ZZZZ) measurements. Also, if the path connects different types of boundaries and has an odd number of kinks, there is a sequence that results in logical CNOT.
@@ 3.3.2 Simplifying Quantum Circuits
These rules can be verified by simple calculations. Note that while the final state must be corrected by the Pauli operation in accordance with the measurement results, the time required for this correction is negligible since this is achieved via the Pauli frame [fowler2012towards, fowler2018low].
@@ 3.4 Solution improvement by 3D routing
This section demonstrates the capabilities and limitations of 3D routing through two contrasting examples. The maximum benefits of 3D routing are observed when the input instruction sequence has a stair-like structure (Fig. 8). For instance, consider an instruction sequence of length $m$ where the $i$-th instruction operates on qubits $i$ and $i+1$. Then, 3D routing achieves a $\Theta(m)$-fold improvement, as it completes the process in $2$ code beats while 2D routing requires $m$ code beats. We emphasize that the subsequent proposed methods aim to replicate this improvement.
@@
Conversely, the benefits vanish when the instruction sequence exhibits a hub-and-spoke structure. Specifically, when the $i$-th instruction connects qubits 11 and $i+1$, the solution obtained by 2D routing is already optimal, meaning 3D routing offers no additional acceleration.
@@ 3.5 Scheduling algorithms
In this section, we propose methods to schedule a sequence of lattice-surgery instructions using the proved theorem. First, we ignore the kink condition for simplicity and discuss four scheduling methods that achieve the 3D routing of lattice-surgery operations. The first two methods, 3D BFS and 3D Dijkstra, are simple extensions of the baseline methods. While they provide scheduling with higher throughput than the baseline method, they need 3D path-finding algorithms and consume longer compilation time. To mitigate this drawback without losing the throughput of instructions, we propose a method named Dijkstra projection. This algorithm provides high-throughput scheduling comparable to 3D variants and achieves reduced time complexity. We also introduce look-ahead Dijkstra projection, which integrates the instruction look-ahead technique to further optimize its scheduling performance. Then, we consider the kink condition and show modifications to Dijkstra projection by which the paths satisfy the kink condition. The baseline and proposed scheduling methods are listed and compared in Tabs. 1 and 2. The time complexities in these tables are analyzed in Sec. 3.6.
Table 1: Comparison of path search subroutines between the scheduling methods. These properties remain the same when instruction look-ahead is integrated. As for the time complexity, $n$ denotes the number of cells in the qubit plane and $h$ denotes the height of the 3D lattice.
@@
Table 2: Comparison of the baseline and proposed scheduling methods. As for the time complexity, $n$ denotes the number of cells in the qubit plane and $m$ denotes the number of instructions.
@@ 3.5.1 Breadth-first search extended in the time direction (3D BFS)
We propose a method called 3D BFS, which utilizes time direction for scheduling with a small modification to BFS. It searches for a path in a 3D lattice, including the time axis, and connects the target data cells by one of the shortest paths. BFS algorithm can be used to find the shortest path in the 3D lattice, as in the case of the 2D lattice. When we choose the endpoint voxels of the lattice-surgery path in the 3D lattice, the choice of their Z coordinates is arbitrary. The 3D BFS method adopts the lowest possible voxel as an endpoint and then searches for the shortest path connecting them. If there is no path connecting them, the height of the endpoint is raised, and the shortest path connecting them is searched for until it is found. For example, by using 3D BFS, the scheduling for the case shown in Fig. 2 is improved to a more efficient result shown in Fig. 10. Note that instruction look-ahead is no longer required to mitigate scheduling delays caused by routing failures, since we can sequentially schedule all the instructions.
Figure 9: Example of 3D BFS.
Figure 10: Example of stacking a path in Dijkstra projection.
The pseudo-code of this procedure is shown in Alg. 3. The dynamic array $A$ maintains whether voxels in the 3D lattice are occupied. The function Find3DShortestPath returns one of the shortest and lowest paths connecting the target data cells, each of which corresponds to $v_{1}$ and $v_{2}$. We can find the path with polynomial time using BFS.

### v4

@@ Abstract
Encoding logical qubits with surface codes and performing multi-qubit logical operations with lattice surgery is one of the most promising approaches to demonstrate fault-tolerant quantum computing. Thus, a method to efficiently schedule a sequence of lattice-surgery operations is vital for high-performance fault-tolerant quantum computing. A possible strategy to improve the throughput of lattice-surgery operations is splitting a large instruction into several small instructions, such as Bell state preparation and measurements, and executing a part of them in advance. However, scheduling methods to fully utilize this idea have yet to be explored. In this paper, we propose a fast and high-performance scheduling algorithm for lattice-surgery instructions leveraging this strategy. We achieved this by converting the scheduling problem of lattice-surgery instructions to a graph problem of embedding 3D paths into a 3D lattice, which enables us to explore efficient scheduling by solving path search problems in the 3D lattice. Based on this reduction, we propose a method to solve the path-finding problems, the look-ahead Dijkstra projection. We numerically show that this method reduced the execution time of benchmark programs generated from quantum phase estimation algorithms by 3.8 times compared with a naive method based on greedy algorithms. Our study establishes the relation between the lattice-surgery scheduling and graph search problems, which leads to further theoretical analysis on compiler optimization of fault-tolerant quantum computing.
@@ 1 Introduction
Quantum computing can efficiently solve several vital scientific problems [38, 23, 28], but its high error rates prevent us from a practical demonstration of quantum advantage. We can overcome this problem with a quantum error-correction technology, i.e., by creating a logical qubit with a negligible error rate by using several noisy qubits. Surface codes [24, 7, 13] are known as promising candidates for quantum error-correcting codes since they show high error-correction performance and can be implemented with qubits having nearest-neighbor interactions on a two-dimensional (2D) grid lattice. Lattice surgery [21, 12] plays a key role in the fault-tolerant quantum computation with surface-code logical qubits: they allow multi-qubit Pauli measurements in a fault-tolerant manner by connecting target logical qubits through ancillary logical-qubit space, and they can be performed in parallel as long as their paths do not overlap. The current state-of-the-art resource estimation stands on architectures with surface codes and lattice surgery [3, 2, 44].
@@
To run quantum algorithms with a short execution time, we should maximize the throughput of quantum instructions [22, 36, 40, 41] by optimizing their scheduling. Unfortunately, finding the optimal scheduling of lattice-surgery instructions is known to be NP-hard [20, 32]. Thus, fast and near-optimal approximation strategies for scheduling lattice-surgery instructions are demanded. For frequently used modules such as magic-state distillation circuits, there are dedicated designs and logical-qubit mappings [12, 26, 27]. On the other hand, for general purposes, we need an algorithm that can accept an arbitrary lattice-surgery sequence. One possible strategy is to map problems into well-known NP-hard instances and find high-quality solutions with fast solvers. Lao et al. [25] map problems into the quadratic assignment problem, and Molavi et al. [32] into SAT problems. While this method would show near-optimal solutions, the target size of the compilation is limited by the capability of solvers. Thus, efficient and high-performance scheduling methods based on other concepts are still in demand.
@@
One of the promising approaches for general and high-performance optimization is to split instructions that are difficult to parallelize into several small fragments and execute parts of them in advance. Beverland et al. [4] proposed a method based on this idea. They propose methods to find a way to split and schedule instructions by finding the edge-disjoint path on a 2D lattice and converting the solution to a lattice-surgery sequence for multiple execution cycles. While this strategy is shown to be effective, the proposed method sticks to specific logical-qubit allocation patterns while possible logical-qubit allocations can be arbitrary [2, 3, 44], which loses the versatility in compilation targets. Also, the range of optimization in this algorithm is limited to instructions within a narrow time window. Therefore, we need a strategy that can fully leverage the idea of instruction split and preceding execution of them.
@@
Here, we propose an efficient and high-performance lattice-surgery scheduling algorithm that utilizes the strategy. The key step of our proposal is establishing the correspondence between a path in a 3D lattice, an executable sequence of lattice-surgery instructions, and achievable actions on target logical qubits. Thanks to this correspondence, we can rephrase the optimization of a lattice-surgery sequence as a problem to embed several 3D paths in a 3D grid lattice, in which we can utilize graph optimization methods. On the basis of this result, we show several polynomial-time approximation algorithms to solve this problem. We evaluated these algorithms with numerical simulations and found that one among them, named look-ahead Dijkstra projection, significantly improves throughput with a reasonable compilation time. We numerically showed that the look-ahead Dijkstra projection improves throughput by 3.8 times compared with a naive solution in a benchmark with realistic applications, SELECT modules in qubitization-based quantum phase estimation [29, 2].
@@
Our results not only optimize the runtime of quantum algorithms but also show that the complicated lattice-surgery compilation can be reduced to simple graph problems, where we can naturally utilize the knowledge of path-search algorithms. Since our algorithm works for general logical-qubit mapping and connectivity, our methods can be extended to other fault-tolerant quantum computing (FTQC) designs based on lattice surgery or distributed quantum computing. While our benchmark focuses on the case of two-body lattice-surgery instructions, we also showed that our method can be straightforwardly extended to multi-body lattice surgery in Appendix A. Thus, our results offer a base for versatile and high-performance compilation for fault-tolerant quantum computing.
@@
  • We propose an efficient and high-performance scheduling method for our 3D formulation, called look-ahead Dijkstra projection. This algorithm leverages the properties of the extended formalism with only a modest increase in time complexity compared to simple baseline algorithms.
@@ 2.1 Surface codes and logical operations
Since qubits suffer from large error rates, we need quantum error correction to suppress their error rate to an arbitrarily small value. Surface codes are known as the most promising quantum error-correcting codes, which encode the information of a qubit into a square cell of qubits. The width of the cell corresponds to a code distance, and by increasing the code distance, we can exponentially suppress the error rates if physical error rates are smaller than the value known as thresholds [24, 14, 1]. A cell of surface codes has two types of boundaries: $X$-type and $Z$-type, where Pauli-$X$ operations along with $X$-type boundary constitute logical $X$ operations, and the same for Pauli-$Z$.
We suppose that qubits are integrated on a large two-dimensional plane, which we call a qubit plane. We assume that the qubit plane is divided into several surface-code logical-qubit cells. During the computation, several cells are used for keeping single-qubit information, which we call data cells. The other cells are available as working space for logical operations on the data cells, which we call ancillary cells. Fig. 1 shows an example of a qubit plane, where dark cells are data cells and light cells are ancillary cells. As surface codes have the two types of boundaries, each cell also has two types of directions: top and bottom are $X$-boundaries, and the others are $Z$-boundaries; or top and bottom are $Z$-boundaries, and the others are $X$-boundaries.
@@ 2.2 Scheduling of lattice surgery
Among the universal operations, the scheduling of lattice-surgery instructions tends to be a complicated task since multiple cells are relevant to them. Thus, in this paper, we focus on the scheduling of lattice-surgery instructions. A lattice-surgery instruction is achieved by connecting target logical-qubit cells. If the target cells are not neighboring, we can use ancillary cells to connect them. Once connected, we need to repeat syndrome measurements for $d$ times, where $d$ is a code distance, to guarantee reliable parity-check measurements. Thus, it is convenient to define this duration of $d$ code cycles for lattice surgery as a time unit, which is called a code beat. The types of connected boundaries determine the basis of Pauli measurements. We can perform Pauli-$Z$ measurements if a $Z$-boundary of a cell is connected and Pauli-$X$ for $X$-boundary.
@@
This example illustrates that optimizing the ordering and routing of lattice surgery is an important factor in maximizing the performance of quantum computing. Unfortunately, the optimal scheduling of the sequence of lattice surgery is known to be NP-hard, i.e., the optimal scheduling is computationally difficult to find within a reasonable time [20]. Thus, an efficient algorithm that can find the near-optimal paths of lattice-surgery instructions is strongly demanded.
@@ 2.3 Baseline scheduling method: BFS
As a baseline method, we introduce a simple procedure to schedule the sequence of lattice-surgery instructions. The most naive strategy for scheduling instructions is to use breadth-first search (BFS) as a subroutine to find one of the shortest paths. We call the procedure BFS, which is defined as follows.
@@
The pseudo-code of this procedure is shown in Alg. 1. Here, we keep whether the ancillary cells are available or not in a map object $A$. FindShortestPath will return one of the shortest paths from $v_{1}$ to $v_{2}$ on graph $G$ for a lattice surgery instruction $o=(v_{1},v_{2},d)$ so that the path does not use surface-code cells marked $1$ in $A$ and $v_{1},v_{2}$ are connected to the path from $d\in\{X,Z\}$ boundaries. If there is no path between them, the function returns an empty list. We can find the shortest paths in a graph $G$ with $V$ vertices and $E$ edges in $O(V+E)$ time using the BFS algorithm. In this procedure, while choosing the shortest path will not necessarily lead to optimal results, we expect this choice to be reasonable since it will consume the fewest cells.
@@ 3.2 Motivating examples
Here, with a simple example, we show that splitting of lattice-surgery instructions enables efficient use of resources, i.e., voxels in the 3D lattice. An overview of our example is shown in Fig. 4. We consider the same situation as Fig. 2. In this situation, the BFS or look-ahead BFS algorithm cannot schedule all the routing at the same code beat, and a hatched path between green cells must wait for the next code beat. The 3D-lattice representation of this scheduling is shown in the left half of the figure. In this case, voxels below green paths are left unused and lose the parallelism of scheduling.
@@
If there is a 3D path that connects XX-($Z$-)boundaries of cells and has an even number of kinks, then there is a sequence of lattice-surgery instructions that consumes the resource corresponding to voxels in the path and results in logical Pauli-XXXX(-ZZZZ) measurements. Also, if the path connects different types of boundaries and has an odd number of kinks, there is a sequence that results in a logical CNOT.
@@ 3.3.2 Simplifying Quantum Circuits
These rules can be verified by simple calculations. Note that while the final state must be corrected by the Pauli operation in accordance with the measurement results, the time required for this correction is negligible since this is achieved via the Pauli frame [14, 12].
@@ 3.4 Solution improvement by 3D routing
This section demonstrates the capabilities and limitations of 3D routing through two contrasting examples. The maximum benefits of 3D routing are observed when the input instruction sequence has a stair-like structure (Fig. 8). For instance, consider an instruction sequence of length $m$ where the $i$-th instruction operates on logical qubits $i$ and $i+1$. Then, 3D routing achieves a $\Theta(m)$-fold improvement, as it completes the process in $2$ code beats while 2D routing requires $m$ code beats. We emphasize that the subsequent proposed methods aim to replicate this improvement.
@@
Conversely, the benefits vanish when the instruction sequence exhibits a hub-and-spoke structure. Specifically, when the $i$-th instruction connects logical qubits 11 and $i+1$, the solution obtained by 2D routing is already optimal, meaning 3D routing offers no additional acceleration.
@@ 3.5 Scheduling algorithms
In this section, we propose methods to schedule a sequence of lattice-surgery instructions using the proved theorem. First, we ignore the kink condition for simplicity and discuss four scheduling methods that achieve the 3D routing of lattice-surgery operations. The first two methods, 3D BFS and 3D Dijkstra, are simple extensions of the baseline methods. While they provide scheduling with higher throughput than the baseline method, they need 3D path-finding algorithms and consume longer compilation time. To mitigate this drawback without losing the throughput of instructions, we propose a method named Dijkstra projection. This algorithm provides high-throughput scheduling comparable to the 3D variants and achieves reduced time complexity. We also introduce the look-ahead Dijkstra projection, which integrates the instruction look-ahead technique to further optimize its scheduling performance. Then, we consider the kink condition and show modifications to the Dijkstra projection algorithms by which the paths satisfy the kink condition. The baseline and proposed scheduling methods are listed and compared in Tabs. 1 and 2. The time complexities in these tables are analyzed in Sec. 3.6.
Table 1: Comparison of path search subroutines between the scheduling methods. These properties remain the same when instruction look-ahead is integrated. As for the time complexity, $n$ denotes the number of cells in the qubit plane, and $h$ denotes the height of the 3D lattice.
@@
Table 2: Comparison of the baseline and proposed scheduling methods. As for the time complexity, $n$ denotes the number of cells in the qubit plane, and $m$ denotes the number of instructions.
@@ 3.5.1 Breadth-first search extended in the time direction (3D BFS)
We propose a method called 3D BFS, which utilizes time direction for scheduling with a small modification to the BFS algorithm. It searches for a path in a 3D lattice, including the time axis, and connects the target data cells by one of the shortest paths. The BFS algorithm can be used to find the shortest path in the 3D lattice, as in the case of the 2D lattice. When we choose the endpoint voxels of the lattice-surgery path in the 3D lattice, the choice of their Z coordinates is arbitrary. The 3D BFS method adopts the lowest possible voxels as endpoints and then searches for the shortest path connecting them. If there is no path connecting them, the height of the endpoint is raised, and the shortest path connecting them is searched for until it is found. For example, by using the 3D BFS, the scheduling for the case shown in Fig. 2 is improved to a more efficient result shown in Fig. 10. Note that instruction look-ahead is no longer required to mitigate scheduling delays caused by routing failures, since we can sequentially schedule all the instructions.
Figure 9: Example of the 3D BFS.
Figure 10: Example of stacking a path in the Dijkstra projection.
The pseudo-code of this procedure is shown in Alg. 3. The dynamic array $A$ maintains whether voxels in the 3D lattice are occupied. The function Find3DShortestPath returns one of the shortest and lowest paths connecting the target data cells, each of which corresponds to $v_{1}$ and $v_{2}$. We can find the path with polynomial time using the BFS.